A student uses a eudiometer tube to collect oxygen gas over water where the temperature is 22 ºC. What is the pressure of the oxygen gas (mm Hg) if the air pressure in the lab is 730 mm Hg?
1
Expert's answer
2018-02-05T07:23:27-0500
In many cases, the amount of gas evolved by a reaction is of interest. Since gases have such small densities, it is usually not practical to collect the gas and find its mass. For gases that are not particularly soluble in water, it is possible to collect the evolved gas by displacement of water from a container.

The setup for the collection of a gas over water involves a container in which the reaction takes place and a gas collection container filled with water and inverted in a reservoir of water. The gas evolved from the reaction is collected by attaching one end of a hose to the reaction container and inserting the other up into the inverted gas collection bottle. As the gas is created, it will displace water from the bottle. The volume of gas can be determined by the amount of water that was displaced by the gas.

The volume of gas collected and the gas laws can be used to calculate the number of moles of gas collected.

During the collection, the water level in the container will adjust so that the pressure inside and outside the container are the same. Because of this, if we know the atmoshperic pressure, we also know the pressure of the gas inside the bottle.

The pressure inside the bottle is partially from the gas being collected and partially from the water vapor that has escaped from the surface of the water in the jar. The water inside the jar will reach an equlibruim state where the number of molecules leaving the surface is the same as the number returning. The equilibrium pressure of water is temperature dependent and is called the vapor pressure of water.

Dalton's Law of Partial Pressures tells us that the total pressure in the container must be the sum of the pressures of the gas we collected and the water vapor.

PT = Pgas + PH2O

T = 22 °C

PH2O = 19.84 mmHg

PT = 730 mmHg

PO2 = PT - PH2O = 730 - 19.84 = 710.16 mmHg